# Black Hole Ringdown Amplitudescopy

### v1

@@ Abstract
Black hole ringdowns in extensions of General Relativity (GR) generically exhibit two distinct signatures: (1) theory-dependent shifts in the standard black-hole quasinormal modes, and (2) additional modes arising from extra fundamental fields – such as scalar, vector, or tensor degrees of freedom – that can also contribute to the gravitational-wave signal. As recently argued, in general both effects are present simultaneously, and accurately modeling them is essential for robust tests of GR in the ringdown regime. In this work, we investigate the impact of extra field-induced modes, which are often neglected in standard ringdown analyses, on the interpretation of gravitational-wave signals. To provide some concrete examples, we focus on dynamical Chern-Simons and Einstein-scalar-Gauss-Bonnet theories, well-motivated extensions of GR, characterized respectively by a parity-odd and a parity-even coupling between a dynamical scalar field and quadratic curvature invariants. We show that including extra field-induced modes improves the bounds on these theories compared to standard spectroscopy and also allows for equally constraining complementary tests not based on quasinormal mode shifts. Our analysis highlights the relevance of incorporating extra field-induced modes in ringdown templates and assesses their potential to either bias or enhance constraints on GR deviations.
@@ I Introduction
Within the framework of linear perturbation theory, the GW signal $h(t)$ at intermediate times after merger is modeled as a sum of damped sinusoids corresponding to the remnant’s QNMs [15], reflecting its relaxation toward a stationary configuration. If the remnant is a BH, GR predicts that its full QNM spectrum is completely determined by only two parameters: the remnant’s mass $M$ and dimensionless spin $\chi$. This distinctive property enables stringent null tests of GR [16, 17, 18, 5, 19, 20, 21, 22, 23] and serves as a robust tool for probing the fundamental nature of the remnant [24, 25, 26]. Very recently, the LIGO-Virgo-KAGRA (LVK) Collaboration performed the first robust spectroscopy test using the exceptional GW250114 event [27, 28], measuring the first overtone with an accuracy of approximately $30\%$, and finding some evidence for other modes in the ringdown of this event.
@@ II Framework
As recently proposed in [40], a general model for the ringdown waveform in theories beyond GR can be written as
@@
In the above expression, $f_{i}^{\rm Kerr}$ and $\tau_{i}^{\rm Kerr}$ (resp., $f_{i}^{{\rm Kerr},\,s=0}$ and $\tau_{i}^{{\rm Kerr},\,s=0}$) are the standard gravitational (resp., scalar) QNM frequency and damping time of a Kerr BH in GR, and depend only on the remnant mass $M$ and spin $\chi$. The quantities $\delta X$’s collectively denote small, theory-dependent deviations from the leading-order quantity $X$. These corrections depend generically on $(M,\chi)$, but also on any fundamental coupling constant of a given GR extension. Their explicit expression must be computed on a case-by-case basis, either perturbatively in the spin [46, 47, 41, 48, 49, 50, 51, 52, 53, 54, 55] or fully numerically [61, 56, 57, 58, 59, 60]. An alternative approach is to parametrize the QNM deviations, either assuming that $\delta X$’s are constant (independent of the BH parameters and theory couplings) [5, 21, 22] or, more realistically, by considering a small-spin expansion of each deviation [62, 63, 64], which inevitably inflates the number of free parameters in the model.
@@ III Examples: Chern-Simons and Gauss-Bonnet gravity theories
The QNMs of spinning BHs in these theories have been originally computed in a small-spin expansion [46, 47, 41, 48, 49, 50, 51, 52, 53, 54, 55] and, more recently, fully numerically [56, 57, 58, 59, 60]. Here we will make use of the fitting formulas for the gravitational modes provided in Table V of Ref. [60] for GB gravity, and Table V of Ref. [68] for CS theory. These fits provide
@@
where $\zeta T_{i}$ is the deviation of the imaginary part of the $i$-th QNM, namely the inverse of the damping time $\tau_{i}$. The dependence on the coupling constant and remnant mass can be factored out into a dimensionless coupling $\zeta=\frac{\alpha^{2}_{\rm CS/GB}}{M^{4}}$, so that the functions $F_{i}$ and $T_{i}$ (for any given $i$-th mode) depend only on the remnant spin, and are of course different for the two theories. In Fig. 1, we plot $F_{i}$ and $T_{i}$ for the most relevant QNMs in these two theories. Note that beyond-GR effects break the isospectrality [70] of the gravitational modes that can be classified as polar and axial in the non-spinning case [48, 71, 72]. As a result, each $(\ell mn)$ mode is further split into an axial (A) and polar (P) mode. As a check, we can observe that, in the zero-spin limit for CS gravity, the deviations of the polar frequencies and damping times approach zero [41].
@@
For a given theory, the normalized amplitudes $\gamma_{\ell mn}$ and the phases $\hat{\phi}_{\ell mn}$ depend on the properties of the progenitor binary. Here we will just consider them as free parameters of the model. Since we will only consider $(220)$ scalar modes, we will simplify the notation by defining $\gamma\equiv\gamma_{220}^{s=0}$.
@@ IV.1 Ringdown Amplitudescopy with O4-like mock data and $\ell=m=2$ modes only
Here, we consider a scenario where a ringdown signal in the modified gravity theories reported above is detected by current-generation interferometers with a representative post-merger signal-to-noise ratio ${\rm SNR}=20$. We assume detection by the LVK network operating at O4 sensitivity, but the specific choice of the sensitivity curve has a mild impact on the analysis. We consider a GW250114-like system [27], with final mass $M=60M_{\odot}$ and final spin $\chi=0.67$. As a reference, the choice of SNR and parameters corresponds to a luminosity distance $d_{L}\approx 236\,{\rm Mpc}$ in O4.
@@
For the results of this section, we assume nearly-symmetric binaries, so that the amplitude of putative higher modes can be neglected [19]. Also, due to the lack of QNM computations and challenges in their modeling [23, 10], we neglect the addition of overtones, although a robust detection of $n=1$ was recently obtained for a similar system [28].
In this first part, we perform an inference test by injecting a signal with AmplPA and recovering it with the same model. We consider the network configurations discussed above and inject a coupling length $\ell_{\rm CS/GB}=35\,{\rm km}$ which, for $M=60M_{\odot}$, corresponds to $\zeta\approx 0.024$. Although this value exceeds current upper bounds [77], it represents a typical scale that can be constrained through BH spectroscopy in representative LVK events [69]. We also inject $A^{R}_{220,A}=0.25$, and do not sample over the $\gamma$ parameter, fixing its value to the injected ones, $\gamma=\gamma_{\rm inj}\in\{0,1,5,10\}$.
@@ IV.2 Ringdown Amplitudescopy with higher modes and future ringdown tests
Here we perform a forecast analysis assuming an event with ringdown ${\rm SNR}=80$ measured by a third-generation interferometer such as the Einstein Telescope (ET) [80, 81, 82, 83]. We optimistically consider a light remnant with $M=20M_{\odot}$, which provides stronger constraints on dimensionful couplings as those arising in quadratic gravity [62]. As a reference, this corresponds to a luminosity distance $d_{L}\approx 437\,{\rm Mpc}$ in ET. We again consider final spin $\chi=0.67$ as a representative value.
@@
In Fig. 3, we show the posterior distributions of the coupling constant $\ell_{\rm GB}$ (left panels) and $\ell_{\rm CS}$ (right panels) for different choices of $A_{R,330}$ and $\gamma_{\rm inj}$, injecting $\ell_{\rm GB/CS}=12\,{\rm km}$. Compared to the previous analysis, we can consider a smaller coupling, as the remnant mass is smaller in this case, corresponding to $\zeta\approx 0.027$. In each panel, we show the results of two analyses: 1) recovery with a model where $\gamma=\gamma_{\rm inj}$ (continuous histograms), and 2) recovery with a model where $\gamma=0$ (dashed histograms), i.e. analog to SpecPA of the previous section but using the $(220)$ and $(330)$ polar modes rather than the $(220)$ axial and polar modes.
The first row of Fig. 3 corresponds to a standard ringdown analysis ($\gamma_{\rm inj}=0$). As expected, when the secondary mode is weakly excited (small $A_{R,330}$), the posterior is almost flat and tends to coincide with the uniform prior $\ell_{\rm GB/CS}\in[0,20]\,{\rm km}$. However, an upper bound at $\ell_{\rm GB}\approx 18$km is still present, because the posterior distribution of $A_{R,330}$ peaks at values higher than the injection, leading to a non-negligible effect of the coupling. When the secondary-mode amplitude is large ($A_{R,330}=0.4$), the injected value of the coupling is well recovered in GB theory, whereas this is not the case for CS theory, since the CS corrections to the $(330)$ polar mode are smaller (see Fig. 1). As a reference, for non-precessing binaries, $A_{R_{,}330}\lesssim 0.5$ for all the mass-ratio and spin configurations [19].
@@
Finally, we also computed the Bayes factors between the biased model and the unbiased model in this case. While the posteriors of the unbiased model shown in Fig. 3 are clearly more accurate for both the GB and the CS cases, we observed that this does not always result in a more discriminating Bayes factor. In the CS case, the Bayes factors grow (in absolute value) monotonically either as $\gamma$ increases or $A_{R,330}$ decreases, as expected. However, in the GB case, the likelihood of the biased distribution is as high as that of the unbiased one, resulting in $\log_{10}{\cal B}\approx 0$ even when $\gamma=10$ and $A_{R,330}=0.005$. This suggests that it would be hard to exclude statistically the biased model, at least based on the Bayes factor estimators. In this case, the maximum likelihood parameters are biased with respect to the injected values, but they still perform a good fit on the model, thus returning an unfaithful Bayes factor, as also observed in other situations [86]. We plan to explore this interesting effect in more detail in future work.
@@ V Conclusions
Given the theory-agnostic character of these modes, it should be straightforward to include them in current LVK pipelines [5], and we advocate for this extension in future ringdown analyses.
As a proof of principle, here we focused on extra scalar modes and non-precessing binaries, but considering other fields (e.g., vectors, tensors), precessing binaries, or modes other than $\ell=m=2,3$ are straightforward extensions [40]. It would also be interesting to consider the effect of extra massive degrees of freedom. Even if the latter do not propagate at relatively low frequencies, they can nevertheless excite the gravitational response and leave an imprint in the ringdown (see [44, 39] for some recent examples).

### v2

@@ Abstract
Black hole ringdowns in extensions of General Relativity (GR) generically exhibit two distinct signatures: i) theory-dependent shifts in the standard black-hole quasinormal modes, and ii) additional modes arising from extra fundamental fields – such as scalar, vector, or tensor degrees of freedom – that can also contribute to the gravitational-wave signal. As recently argued, in general both effects are present simultaneously, and accurately modeling them is essential for robust tests of GR in the ringdown regime. In this work, we investigate the impact of extra field-induced modes, which are often neglected in standard ringdown analyses, on the interpretation of gravitational-wave signals. To provide some concrete examples, we focus on dynamical Chern-Simons and Einstein-scalar-Gauss-Bonnet theories, well-motivated extensions of GR, characterized respectively by a parity-odd and a parity-even coupling between a dynamical scalar field and quadratic curvature invariants. We show that including extra field-induced modes improves the bounds on these theories compared to standard spectroscopy and also allows for equally constraining complementary tests not based on quasinormal mode shifts. Our analysis highlights the relevance of incorporating extra field-induced modes in ringdown templates and assesses their potential to either bias or enhance constraints on GR deviations.
@@ I Introduction
Within the framework of linear perturbation theory, the GW signal $h(t)$ at intermediate times after merger is modeled as a sum of damped sinusoids corresponding to the remnant’s QNMs [75], which reflects the BH relaxation toward a stationary configuration. If the remnant is a BH, GR predicts that its full QNM spectrum is completely determined by only two parameters: the remnant’s mass $M$ and dimensionless spin $\chi$. This distinctive property enables stringent null tests of GR [26, 32, 68, 5, 55, 56, 80, 79, 10] and serves as a robust tool for probing the fundamental nature of the remnant [81, 82, 83]. Very recently, the LIGO-Virgo-KAGRA (LVK) Collaboration performed the first robust spectroscopy test using the exceptional GW250114 event [3, 20], constraining the first overtone frequency with an accuracy of approximately $30\%$, and finding some evidence for other modes in the ringdown of this event.
@@ II Framework
As recently proposed in [45], a general model for the ringdown waveform in theories beyond GR can be written as  [^1] For simplicity, we assume circular polarization.
@@
In the above expression, $f_{i}^{\rm Kerr}$ and $\tau_{i}^{\rm Kerr}$ (resp., $f_{i}^{{\rm Kerr},\,s=0}$ and $\tau_{i}^{{\rm Kerr},\,s=0}$) are the standard gravitational (resp., scalar) QNM frequency and damping time of a Kerr BH in GR, and depend only on the remnant mass $M$ and spin $\chi$. The quantities $\delta X$’s collectively denote small, theory-dependent deviations from the leading-order quantity $X$. These corrections depend generically on $(M,\chi)$, but also on any fundamental coupling constant of a given GR extension. Their explicit expression must be compute on a case-by-case basis, either perturbatively in the spin [92, 29, 87, 90, 91, 95, 103, 27, 96, 28, 102] or fully numerically [48, 38, 39, 22, 41, 40]. An alternative approach is to parametrize the QNM deviations, either assuming that $\delta X$’s are constant (independent of the BH parameters and theory couplings) [5, 80, 79] or, more realistically, by considering a small-spin expansion of each deviation [85, 33, 86], which inevitably inflates the number of free parameters in the model.
@@ III Examples: Chern-Simons and Gauss-Bonnet gravity theories
The QNMs of spinning BHs in these theories have been originally compute in a small-spin expansion [92, 29, 87, 90, 91, 95, 103, 27, 96, 28, 102] and, more recently, fully numerically [38, 39, 22, 41, 40]. Here we will make use of the fitting formulas for the gravitational modes provided in Table V of Ref. [40] for GB gravity, and Table V of Ref. [37] for CS theory. These fits provide
@@
where $\zeta T_{i}$ is the deviation of the imaginary part of the $i$-th QNM, namely the inverse of the damping time $\tau_{i}$. In Appendix A, we review a derivation of Eq. (7) and provide the fitting coefficients. The dependence on the coupling constant and remnant mass can be factored out into a dimensionless coupling $\zeta=\frac{\alpha^{2}_{\rm CS/GB}}{M^{4}}$, so that the functions $F_{i}$ and $T_{i}$ (for any given $i$-th mode) depend only on the remnant spin, and are of course different for the two theories. In Fig. 1, we plot $F_{i}$ and $T_{i}$ for the most relevant QNMs in these two theories. Note that beyond-GR effects break the isospectrality [35] of the gravitational modes that can be classified as polar and axial in the non-spinning case [90, 12, 77]. As a result, each $(\ell mn)$ mode is further split into an axial (A) and polar (P) mode. As a check, we can observe that, in the zero-spin limit for CS gravity, the deviations of the polar frequencies and damping times approach zero [87].
@@
Putative higher-order corrections are consistently neglected within our framework. For a given theory, the normalized amplitudes $\gamma_{\ell mn}$ and the phases $\hat{\phi}_{\ell mn}$ depend on the properties of the progenitor binary. Here we will just consider them as free parameters of the model. Since we will only consider $(220)$ scalar modes, we will simplify the notation by defining $\gamma\equiv\gamma_{220}^{s=0}$.
@@ IV.1 Ringdown Amplitudescopy with O4-like mock data and $\ell=m=2$ modes only
Here, we consider a scenario where a ringdown signal in the modified gravity theories reported above is detected by current-generation interferometers with a representative ringdown optimal signal-to-noise ratio of ${\rm SNR}=20$. We assume detection by the LVK network operating at O4 sensitivity provided in Ref. [78], but the specific choice of the sensitivity curve has a mild impact on the analysis. We consider a GW250114-like system [3], with final mass $M=60M_{\odot}$ and final spin $\chi=0.67$. As a reference, the choice of SNR (computed starting at the peak of the ringdown signal) and parameters corresponds to a luminosity distance $d_{L}\approx 236\,{\rm Mpc}$ in O4.
@@
For the results of this section, we assume nearly-symmetric binaries, so that the amplitude of putative odd higher modes can be neglected [55]. Furthermore, we neglect the possible effect of $\ell=m=4$ modes. Although some evidence of the latter was identified in GW250114, they were poorly measured [20]. Also, due to the lack of QNM computations and challenges in their modeling [10, 16], we neglect the addition of overtones, although a robust detection of $n=1$ was recently obtained for GW250114 [20].
In this first part, we perform an inference test by injecting a signal with AmplPA and recovering it with the same model. We consider the network configurations discussed above and inject a coupling length $\ell_{\rm CS/GB}=35\,{\rm km}$ which, for $M=60M_{\odot}$, corresponds to $\zeta\approx 0.024$. Although this value exceeds current upper bounds [97], it represents a typical scale that can be constrained through BH spectroscopy in representative LVK events [42]. We also inject $A^{R}_{220,A}=0.25$, and do not sample over the $\gamma$ parameter, fixing its value to the injected ones, $\gamma=\gamma_{\rm inj}\in\{0,1,5,10\}$. We chose to fix $\gamma$ for two reasons. The first is because, ideally, for a given theory, one could have an estimate of that value in terms of the progenitor parameters from simulations, which are finally available for certain theories. The second is more technical: indeed, $\gamma$ is degenerate with final mass and $\alpha$, and its measurement might be difficult. Nonetheless, in Appendix B we relax this assumption and include $\gamma$ in the sampling parameters, obtaining similar results as those presented here.
@@
We also have performed the analyses of Fig. 2 by sampling with a uniform prior on $\gamma$, and we report the results in Appendix B for the case of CS, showing that sampling on $\gamma$ has a small impact on the results.
@@ IV.2 Ringdown Amplitudescopy with higher modes and future ringdown tests
Here we perform a forecast analysis assuming an event with ringdown ${\rm SNR}=80$ measured by a third-generation interferometer such as the Einstein Telescope (ET) [62, 84, 24, 4]. We optimistically consider a light remnant with $M=20M_{\odot}$, which provides stronger constraints on dimensionful couplings as those arising in quadratic gravity [85]. As a reference, this corresponds to a luminosity distance $d_{L}\approx 437\,{\rm Mpc}$ in ET. We again consider the final spin $\chi=0.67$ as a representative value. We assume ET in a triangle configuration with arm length of 15km, with the PSD provided in Ref. [24]. Also in this case, the impact of the different PSDs has a mild impact on the analysis, and the impact of the BH spectroscopy results due to the choice of the detector configuration is discussed in Ref. [24].
@@
In Fig. 3, we show the posterior distributions of the coupling constant $\ell_{\rm GB}$ (left panels) and $\ell_{\rm CS}$ (right panels) for different choices of $A_{R,330}$  [^4] Notice that the values of $A_{\rm R,330}$ reach values lower than 0.04, corresponding to mass ratio very close to unity. Even if these values cannot be considered as a confident detection of the 330 mode, we report it anyway, as we can analyze the impact of adding a third mode with different excitations. and $\gamma_{\rm inj}$, injecting $\ell_{\rm GB/CS}=12\,{\rm km}$, while the full corner plots are reported in Appendix B. Compared to the previous analysis, we can consider a smaller coupling, as the remnant mass is smaller in this case, corresponding to $\zeta\approx 0.027$. In each panel, we show the results of two analyses: 1) recovery with a model where $\gamma=\gamma_{\rm inj}$ (continuous histograms), and 2) recovery with a model where $\gamma=0$ (dashed histograms), i.e. analog to SpecPA of the previous section but using the $(220)$ and $(330)$ polar modes rather than the $(220)$ axial and polar modes.
The first row of Fig. 3 corresponds to a standard ringdown analysis ($\gamma_{\rm inj}=0$). As expected, when the secondary mode is weakly excited (small $A_{R,330}$), the posterior is almost flat and tends to coincide with the uniform prior $\ell_{\rm GB}\in[0,16]\,{\rm km}$ and $\ell_{\rm CS}\in[0,20]\,{\rm km}$, respectively  [^5] We did not consider to higher values of $\ell_{\rm GB}$ due to stability issue of the simulations related to the frequency shifts reported in Ref. [40]. . However, we find some interesting behaviors for both the theories. For GB, we find that the distribution in $\ell_{\rm GB}$ shifts near the upper bound of the prior for low $A_{330,R}$ values. We suspect that this behavior is given by a secondary peak in the final mass distribution (see Appendix B), which might be due to a poor convergence of the analysis when the 330 mode has a low amplitude. Indeed, since the beyond-GR effects depend on the combination $\zeta=\alpha^{2}/M^{4}$, a bias in the mass correlates with a bias in the coupling. However, this bias is not an actual problem, given the low value of the injected $A_{R,330}$ and the fact that the recovery is correct for more realistic values of $A_{R,330}$. For CS instead, an upper bound at $\ell_{\rm CS}\approx 18$km is present, because the posterior distribution of $A_{R,330}$ peaks at values higher than the injection, leading to a non-negligible effect of the coupling. When the secondary-mode amplitude is large ($A_{R,330}=0.4$), the injected value of the coupling is well recovered in GB theory, whereas this is not the case for CS theory, since the CS corrections to the $(330)$ polar mode are smaller (see Fig. 1). As a reference, for non-precessing binaries, $A_{R_{,}330}\lesssim 0.5$ for all the mass-ratio and spin configurations [55].
@@
Finally, we also compute the Bayes factors between the biased model and the unbiased model in this case. While the posteriors of the unbiased model shown in Fig. 3 are clearly more accurate for both the GB and the CS cases, we observe that this does not always result in a more discriminating Bayes factor. In the CS case, the Bayes factors grow (in absolute value) monotonically either as $\gamma$ increases or $A_{R,330}$ decreases, as expected. However, in the GB case, the likelihood of the biased distribution is as high as that of the unbiased one, resulting in $\log_{10}{\cal B}\approx 0$ even when $\gamma=10$ and $A_{R,330}=0.005$. This suggests that it would be hard to exclude statistically the biased model, at least based on the Bayes factor estimators. In this case, the maximum likelihood parameters are biased with respect to the injected values, but they still perform a good fit on the model, thus returning an unfaithful Bayes factor, as also observed in other situations [67]. This feature for GB theory might be due to the fact that the posterior of $\ell_{GB}$ is constrained near the prior upper bound, which is close to the injected value. As a future extension, it might be interesting to understand if the small values of $\log_{10}(\mathcal{B})$ persist when adding further modes in addition to 330P. For instance, it would be interesting to merge the analyses performed in Sections IV.1 and IV.2, thus including 220P+220A+330P plus, possibly, a scalar mode. More in general, an accurate forecast should be performed, to understand which are the modes that impact the most in the analysis and which ones can mask the scalar contribution.
@@ V Conclusions
In this work, since we have injected a ringdown waveform, we are neglecting the issue of the amplitude stability [36], as the true signal does not contain nonlinearities and transient effects. However, when performing real data analysis, the amplitude stability is reached only at late times. In this last case, there can be two ways of performing the amplitudescopy analysis: either starting at a fixed time sufficiently far after the merger (see, for instance, in [42]), or improving currently beyond-GR IMR waveform models (e.g. [71]). We leave these improvements for future works.
Further, as a proof of principle, here we focused on extra scalar modes and non-precessing binaries, but considering other fields (e.g., vectors, tensors), precessing binaries, or modes other than $\ell=m=2,3$ are straightforward extensions [45]. It would also be interesting to consider the effect of extra massive degrees of freedom. Even if the latter do not propagate at relatively low frequencies, they can nevertheless excite the gravitational response and leave an imprint in the ringdown (see [30, 8] for some recent examples).